THE FORMAMIDE-(H_{2}O)_{n} (n=1,2) CLUSTERS BY ROTATIONAL SPECTROSCOPY

SUSANA BLANCO, ALBERTO LESARRI, \underline{JUAN C. LÓPEZ} and JOSÉ L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain.

Formamide (NH_{2}-COH) is the most simple molecule carrying the peptide linkage and its microsolvated clusters can be used to investigate the diverse hydrogen bond interactions of water with this functional group. A molecular beam Fourier Transform microwave spectrometer^{a} equipped with a heating nozzle^{b} has been used for this research. The most stable conformer of the formamide-H_{2}O complex, previously observed by Lovas et al.,^{c} is stabilized by two hydrogen bonds O-H⋯O=C and N-H⋯O with water closing a cycle with the peptide functional group. In the present work we have observed two new formamide-H_{2}O conformers and the trimer formamide-(H_{2}O)_{2}. In the second most stable 1:1 conformer water is placed on the other side of the carbonyl group stabilized by a O-H⋯O=C hydrogen bond and by a C-H⋯O weak hydrogen bond. In the third 1:1 complex, water is bonded to the amino group by a N-H⋯O hydrogen bond. For the formamide-(H_{2}O)_{2} complex the two molecules of water close a cycle with the peptide group with three different hydrogen bonds (O-H⋯O=C, O-H⋯O and N-H⋯O). The rotational spectra of the parent and several D/H, ^{15}N/^{14}N, ^{13}C/^{12}C and ^{18}O/^{16}O isotopically substituted species have been measured for the formamide monomer, the most stable conformer of the formamide-H_{2}O adduct and the formamide-(H_{2}O)_{2} complex in order to investigate their structures. High level ab initio computations have been performed in order to complement the experimental results.

a J. L. Alonso, F. Lorenzo, J. C. López, A. Lesarri, S. Mata and H. Dreizler; Chem. Phys., 218, 267 (1997)

b S. Blanco, J. C. López, J. L. Alonso, P. Ottaviani and W. Caminati; J. Chem. Phys., 119, 880 (2003)

c F. J. Lovas, R. D. Suenram, G. T. Fraser, C. W. Gillies and J. Zozom, J. Chem. Phys., 88 722 (1988)